[Ian, Torrey]
Procedure notes in 775 cavity progress in the lab today:
-Improved alignment in vertical direction. Alignment is sufficient for initial tests and cavity characterization but can be improved. Mode scan suggests that mode matching could be improved.
-Polarization fluctuations in fiber due to fans is problematic because of how we are attenuating the power (PBS). We have reduced the overall laser power now that alignment is done until 1811 PD in not saturated. This is done by turning the amplifier all the way down and removing the PBS. The 1811 is still saturated in this configuration. Temporarily changed the SHG polarization so less power goes into fiber on the SHG sled to the OFC sled.
-This fixed any problems with the refl PD we now have a robust lock (with the laser).
-Peak to peak error amplitude with no tank circuit and 5 mV set point offset is between 10.5 mV to 11 mV. Tank circuit was stolen from the 1550nm EOM as the capacitance of the 775nm EOM is the same (14 pF). (Note that we have the parts to build another LC circuit on the 1550nm path but are waiting for the little circuit boxes that should arrive with the big thorlabs order that has been placed.) Error amp is now 34 mV requiring an error set point of 11 mV. We did not adjust the modulation frequency of 50.68 mHz, could potentially get some more amplitude from this.
-The cavity was then locked with the laser and we took measurements for G and the noise spectrum at the error point in the same way as previous posts. Data for this can be found in the nextcloud: "\Nextcloud\GQuEST\B102\Output Filter Cavity\775nm locks\Initial laser lock"